# The numeraire e-variable and reverse information projection

### v3

@@ 3 Duality and reverse information projection
(i) $\Leftrightarrow$ (ii): Suppose ${\mathsf{P}}^{*}$ is the RIPr. Because ${\mathsf{P}}^{*}$ and ${\mathsf{Q}}$ are equivalent we have $d{\mathsf{P}}^{a}/d{\mathsf{P}}^{*}=(d{\mathsf{P}}^{a}/d{\mathsf{Q}})/(d{\mathsf{P}}^{*}/d{\mathsf{Q}})=X^{*}d{\mathsf{P}}^{a}/d{\mathsf{Q}}$ up to ${\mathsf{Q}}$-nullsets for any ${\mathsf{P}}\in{\mathcal{P}}^{\circ\circ}$. Thus we obtain (ii) by taking $X=X^{*}$ in (3.2). This proves the forward direction. To prove the converse, we show that the RIPr is the unique element of ${\mathcal{P}}^{\circ\circ}$ that is equivalent to ${\mathsf{Q}}$ and satisfies (ii). This is argued exactly as uniqueness of the numeraire: if ${\mathsf{P}}_{1}^{*},{\mathsf{P}}_{2}^{*}\in{\mathcal{P}}^{\circ\circ}$ are both equivalent to ${\mathsf{Q}}$ and satisfy ${\mathbb{E}}_{\mathsf{Q}}[{d{\mathsf{P}}^{a}}/{d{\mathsf{P}}_{i}^{*}}]\leq 1$ for all ${\mathsf{P}}\in{\mathcal{P}}^{\circ\circ}$, then with $Y=d{\mathsf{P}}^{*}_{2}/d{\mathsf{P}}^{*}_{1}$ we have $1\leq 1/{\mathbb{E}}_{\mathsf{Q}}[Y]\leq{\mathbb{E}}_{\mathsf{Q}}[1/Y]\leq 1$. Thus $Y$ is ${\mathsf{Q}}$-almost surely equal to one and ${\mathsf{P}}^{*}_{1}={\mathsf{P}}^{*}_{2}$.
(ii) $\Leftrightarrow$ (iii): The forward implication follows from Jensen’s inequality, and the converse uses the same argument as the proof of Proposition 2.4. Specifically, ${\mathsf{P}}(t)=(1-t){\mathsf{P}}^{*}+t{\mathsf{P}}$ is in ${\mathcal{P}}^{\circ\circ}$ for all $t\in(0,1)$, so ${\mathbb{E}}_{\mathsf{Q}}[t^{-1}\log(d{\mathsf{P}}(t)^{a}/d{\mathsf{P}}^{*})]\leq 0$. Sending $t$ to zero yields (ii).
@@
for all $X\in{\mathcal{E}}$ and ${\mathsf{P}}\in{\mathcal{P}}^{\circ\circ}$. This implies the ‘weak duality’ inequality
@@
The strong duality in (3.3) implies that there exists an e-variable with positive expected logarithm under ${\mathsf{Q}}$ if and only if ${\mathsf{Q}}\notin{\mathcal{P}}^{\circ\circ}$. Therefore, thanks to Corollary 4.6 below, Theorem 3.4 extends a result by Zhang et al. (2023) showing that if ${\mathcal{P}}$ is finite, then such an e-variable exists if and only if ${\mathsf{Q}}$ is not in the convex hull of ${\mathcal{P}}$.
@@
The RIPr ${\mathsf{P}}^{*}$ is not a probability measure in general. Equivalently, it may happen that ${\mathbb{E}}_{\mathsf{Q}}[1/X^{*}]<1$. We will see this phenomenon in Sections 5.1 and 5.4, and other examples are given by Lardy et al. (2023). It is also worth noting that there may exist elements $\widetilde{\mathsf{P}}\in{\mathcal{P}}^{\circ\circ}$ that are different from ${\mathsf{P}}^{*}$, but whose absolutely continuous part $\widetilde{\mathsf{P}}^{a}$ is equal to ${\mathsf{P}}^{*}$. We will see this in Section 5.4 with $\widetilde{\mathsf{P}}$ being the symmetrization of ${\mathsf{Q}}$. Other examples are obtained from Example 3.7 below.
@@
Consider a simple null hypothesis ${\mathcal{P}}=\{{\mathsf{P}}_{0}\}$ with ${\mathsf{Q}}\ll{\mathsf{P}}_{0}$ so that the numeraire $X^{*}$ is finite. In this case ${\mathsf{P}}^{*}={\mathsf{P}}_{0}^{a}$ and $X^{*}=1/(d{\mathsf{P}}^{*}/d{\mathsf{Q}})=d{\mathsf{Q}}/d{\mathsf{P}}_{0}$ up to ${\mathsf{Q}}$-nullsets. Here one can easily check that ${\mathcal{P}}^{\circ\circ}$ consists of all sub-probability measures ${\mathsf{P}}$ such that ${\mathsf{P}}(A)\leq{\mathsf{P}}_{0}(A)$ for all $A$. Indeed, any such ${\mathsf{P}}$ must belong to ${\mathcal{P}}^{\circ\circ}$ since ${\mathbb{E}}_{\mathsf{P}}[X]\leq{\mathbb{E}}_{{\mathsf{P}}_{0}}[X]\leq 1$ for all e-variables $X$. Conversely, if ${\mathsf{P}}$ is not of this form, then ${\mathsf{P}}(A)>{\mathsf{P}}_{0}(A)$ for some $A$, and $X=\bm{1}_{A}/{\mathsf{P}}_{0}(A)$ is an e-variable such that ${\mathbb{E}}_{\mathsf{P}}[X]>1$, showing that ${\mathsf{P}}\notin{\mathcal{P}}^{\circ\circ}$. If ${\mathsf{P}}_{0}(A)=0$ we set $X=\infty\bm{1}_{A}$ in the previous sentence.
We end this section with the general version of Theorem 3.4. The proof reduces the general case to the one treated in Theorem 3.4 by introducing the conditional probability measure ${\mathsf{Q}}^{*}(A)={\mathsf{Q}}(A\mid X^{*}<\infty)$ if the quantity $\lambda^{*}={\mathsf{Q}}(X^{*}<\infty)$ is positive, where $X^{*}$ is the numeraire. The fully degenerate case $\lambda^{*}=0$ is treated separately. It is evident from Definition 3.2 that the RIPr ${\mathsf{P}}^{*}$ is absolutely continuous with respect to ${\mathsf{Q}}$, and we know from Lemma 3.1 that it belongs to ${\mathcal{P}}^{\circ\circ}$. As shown below, one actually has more precise information, namely that
@@
where we write $\lambda^{*}\cdot{\mathcal{P}}^{\circ\circ}$ for the set $\{\lambda^{*}{\mathsf{P}}\colon{\mathsf{P}}\in{\mathcal{P}}^{\circ\circ}\}\subset{\mathcal{P}}^{\circ\circ}$. We continue to write ${\mathsf{P}}^{a}$ for the absolutely continuous part of ${\mathsf{P}}$ with respect to ${\mathsf{Q}}$.
@@
    ${\mathbb{E}}_{\mathsf{Q}}\left[\frac{d{\mathsf{P}}^{a}}{d{\mathsf{P}}^{*}}\bm{1}_{\{X^{*}<\infty\}}\right]\leq 1\text{ for all }{\mathsf{P}}\in{\mathcal{P}}^{\circ\circ}.$
@@
    ${\mathbb{E}}_{\mathsf{Q}}\left[\log\left(\lambda^{*}\frac{d{\mathsf{P}}^{a}}{d{\mathsf{P}}^{*}}\right)\bm{1}_{\{X^{*}<\infty\}}\right]\leq 0\text{ for all }{\mathsf{P}}\in{\mathcal{P}}^{\circ\circ}.$
@@
${\mathbb{E}}_{{\mathsf{Q}}^{*}}[\log X^{*}]=\sup_{X\in{\mathcal{E}}}{\mathbb{E}}_{{\mathsf{Q}}^{*}}[\log X]=\inf_{{\mathsf{P}}\in{\mathcal{P}}^{\circ\circ}}H({\mathsf{Q}}^{*}\mid{\mathsf{P}})=H({\mathsf{Q}}^{*}\mid{\mathsf{P}}^{**}).$ (3.7)
@@
We first let ${\mathsf{P}}^{*}$ be the RIPr and check (3.6). Equivalence with ${\mathsf{Q}}$ on $\{X^{*}<\infty\}$ is evident from Definition 3.2. To show that ${\mathsf{P}}^{*}$ belongs to $\lambda^{*}\cdot{\mathcal{P}}^{\circ\circ}$ we refine the calculation in the proof of Lemma 3.1. Consider any $X\in{\mathcal{E}}$ and set $X^{\prime}=X\bm{1}_{\{X^{*}<\infty\}}+\infty\bm{1}_{\{X^{*}=\infty\}}$. This still belongs to ${\mathcal{E}}$ since it is ${\mathsf{P}}$-almost surely equal to $X$ for all ${\mathsf{P}}\in{\mathcal{P}}$. Thanks to the convention (1.1) we have $X^{\prime}/X^{*}=\bm{1}_{\{X^{*}<\infty\}}X^{\prime}/X^{*}+\bm{1}_{\{X^{*}=\infty\}}$ and hence
@@
using the numeraire property in the last step. It follows that ${\mathsf{P}}^{*}\in\lambda^{*}\cdot{\mathcal{P}}^{\circ\circ}$.
@@
Since ${\mathsf{Q}}^{*}(X^{*}<\infty)=1$ by construction, we have from Theorem 2.5 that ${\mathsf{Q}}^{*}\ll{\mathcal{P}}$. We also define ${\mathsf{P}}^{**}={\mathsf{P}}^{*}/\lambda^{*}$ which, thanks to (3.6), is an element of ${\mathcal{P}}^{\circ\circ}$ equivalent to ${\mathsf{Q}}^{*}$. Now, the statement (i) that ${\mathsf{P}}^{*}$ is the RIPr of ${\mathsf{Q}}$ is equivalent to ${\mathsf{P}}^{**}$ being the RIPr of ${\mathsf{Q}}^{*}$. Furthermore, (ii) and (iii) can be equivalently expressed in terms of ${\mathsf{Q}}^{*}$ and ${\mathsf{P}}^{**}$ as
${\mathbb{E}}_{{\mathsf{Q}}^{*}}\left[\frac{d{\mathsf{P}}^{a}}{d{\mathsf{P}}^{**}}\right]\leq 1\text{ for all }{\mathsf{P}}\in{\mathcal{P}}^{\circ\circ}\quad\text{and}\quad{\mathbb{E}}_{{\mathsf{Q}}^{*}}\left[\log\frac{d{\mathsf{P}}^{a}}{d{\mathsf{P}}^{**}}\right]\leq 0\text{ for all }{\mathsf{P}}\in{\mathcal{P}}^{\circ\circ},$
@@
If ${\mathsf{P}}_{0}$ and ${\mathsf{Q}}$ are singular we have ${\mathsf{P}}^{*}=0$ and $X^{*}=\infty$ which are indeed the RIPr and numeraire, so we focus on the non-singular case. We first compute the numeraire. Thanks to Example 3.7, the numeraire for ${\mathsf{Q}}^{*}$ is $d{\mathsf{Q}}^{*}/d{\mathsf{P}}_{0}$ up to ${\mathsf{Q}}^{*}$-nullsets. The numeraire for ${\mathsf{Q}}$ should be infinite on $N$, which leads to (3.8). We note that $X^{*}$ is ${\mathsf{Q}}$-almost surely strictly positive, it is an e-variable because ${\mathbb{E}}_{{\mathsf{P}}_{0}}[X^{*}]=1$, and for any other e-variable $X$ we have ${\mathbb{E}}_{\mathsf{Q}}[X/X^{*}]\leq{\mathsf{Q}}(N)+{\mathsf{Q}}(N^{c}){\mathbb{E}}_{{\mathsf{Q}}^{*}}[X/(d{\mathsf{Q}}^{*}/d{\mathsf{P}}_{0})]\leq 1$. This confirms that $X^{*}$ is the numeraire, and since $1/X^{*}=d{\mathsf{P}}_{0}^{a}/d{\mathsf{Q}}$ up to ${\mathsf{Q}}$-nullsets the RIPr is ${\mathsf{P}}^{*}={\mathsf{P}}_{0}^{a}$ as claimed. ∎
4 Finding the numeraire
So far we have studied the abstract existence, uniqueness, and duality theory of the numeraire, but we have not yet said much about how to actually compute it. In this section we develop tools for doing so, which we then apply to several examples in Section 5. We focus on the statistically most relevant situation where ${\mathsf{Q}}\ll{\mathcal{P}}$. In this case the numeraire $X^{*}$, and indeed every e-variable, is ${\mathsf{Q}}$-almost surely finite and the RIPr ${\mathsf{P}}^{*}$ is equivalent to ${\mathsf{Q}}$. The general case can be reduced to this case by replacing ${\mathsf{Q}}$ with the conditional measure ${\mathsf{Q}}^{*}$ in Theorem 3.8 (except in the fully degenerate situation where the numeraire is infinity and the RIPr zero.)
Proposition 2.4 suggests that $X^{*}$ can be found by solving an optimization problem. Unfortunately this may not be straightforward because the optimization is over the set of all e-variables, which can be difficult to characterize. For the same reason, even if a candidate $X^{*}$ has been found it may not be clear how to check that it does, in fact, satisfy the numeraire property. Alternatively, the strong duality of Theorem 3.4 suggests minimizing relative entropy to find ${\mathsf{P}}^{*}$. However, the optimization is again over a set that can be difficult to describe explicitly, in this case the effective null ${\mathcal{P}}^{\circ\circ}$. Moreover, checking that a candidate ${\mathsf{P}}^{*}$ is optimal involves comparing it to all elements of ${\mathcal{P}}^{\circ\circ}$.
We first present a verification theorem which simplifies the task of checking that candidates $X^{*}$ and ${\mathsf{P}}^{*}$ are in fact optimal. Despite the simple proof, the result turns out to be remarkably useful.

### v4

@@ 3 Duality and reverse information projection
(i) $\Leftrightarrow$ (ii): Suppose ${\mathsf{P}}^{*}$ is the RIPr. Because ${\mathsf{P}}^{*}$ and ${\mathsf{Q}}$ are equivalent we have $d{\mathsf{P}}^{a}/d{\mathsf{P}}^{*}=(d{\mathsf{P}}^{a}/d{\mathsf{Q}})/(d{\mathsf{P}}^{*}/d{\mathsf{Q}})=X^{*}d{\mathsf{P}}^{a}/d{\mathsf{Q}}$ up to ${\mathsf{Q}}$-nullsets for any ${\mathsf{P}}\in{\mathcal{P}_{\textnormal{eff}}}$. Thus we obtain (ii) by taking $X=X^{*}$ in (3.2). This proves the forward direction. To prove the converse, we show that the RIPr is the unique element of ${\mathcal{P}_{\textnormal{eff}}}$ that is equivalent to ${\mathsf{Q}}$ and satisfies (ii). This is argued exactly as uniqueness of the numeraire: if ${\mathsf{P}}_{1}^{*},{\mathsf{P}}_{2}^{*}\in{\mathcal{P}_{\textnormal{eff}}}$ are both equivalent to ${\mathsf{Q}}$ and satisfy ${\mathbb{E}}_{\mathsf{Q}}[{d{\mathsf{P}}^{a}}/{d{\mathsf{P}}_{i}^{*}}]\leq 1$ for all ${\mathsf{P}}\in{\mathcal{P}_{\textnormal{eff}}}$, then with $Y=d{\mathsf{P}}^{*}_{2}/d{\mathsf{P}}^{*}_{1}$ we have $1\leq 1/{\mathbb{E}}_{\mathsf{Q}}[Y]\leq{\mathbb{E}}_{\mathsf{Q}}[1/Y]\leq 1$. Thus $Y$ is ${\mathsf{Q}}$-almost surely equal to one and ${\mathsf{P}}^{*}_{1}={\mathsf{P}}^{*}_{2}$.
(ii) $\Leftrightarrow$ (iii): The forward implication follows from Jensen’s inequality, and the converse uses the same argument as the proof of Proposition 2.4. Specifically, ${\mathsf{P}}(t)=(1-t){\mathsf{P}}^{*}+t{\mathsf{P}}$ is in ${\mathcal{P}_{\textnormal{eff}}}$ for all $t\in(0,1)$, so ${\mathbb{E}}_{\mathsf{Q}}[t^{-1}\log(d{\mathsf{P}}(t)^{a}/d{\mathsf{P}}^{*})]\leq 0$. Sending $t$ to zero yields (ii).
@@
for all $X\in{\mathcal{E}}$ and ${\mathsf{P}}\in{\mathcal{P}_{\textnormal{eff}}}$. This implies the ‘weak duality’ inequality
@@
The equivalence of (ii) and (iii) in Theorem 3.6, and the equivalence stated in Proposition 2.4, have commonalities that can be understood by considering two optimization problems. The first is to maximize $f(X)={\mathbb{E}}_{\mathsf{Q}}[\log(X/X^{*})]$ over $X\in{\mathcal{E}}$. The second is to minimize $g({\mathsf{P}})={\mathbb{E}}_{\mathsf{Q}}[-\log(d{\mathsf{P}}^{a}/d{\mathsf{P}}^{*})]$ over ${\mathsf{P}}\in{\mathcal{P}_{\textnormal{eff}}}$. The numeraire property can be interpreted as the first-order condition that the Gateaux (i.e. directional) derivative $f^{\prime}(X^{*};X-X^{*})=\lim_{t\downarrow 0}t^{-1}(f(X^{*}+t(X^{*}-X))-f(X^{*}))$ in any feasible direction is nonpositive at $X^{*}$. Thanks to concavity, this is necessary and sufficient for optimality. Analogously, property (ii) of Theorem 3.6 states that the Gateaux derivative $g^{\prime}({\mathsf{P}}^{*};{\mathsf{P}}-{\mathsf{P}}^{*})=\lim_{t\downarrow 0}t^{-1}(g({\mathsf{P}}^{*}+t({\mathsf{P}}^{*}-{\mathsf{P}}))-g({\mathsf{P}}^{*}))$ in any feasible direction is nonnegative at ${\mathsf{P}}^{*}$. This is again necessary and sufficient for optimality, expressed by (iii). The first-order condition for $f$ states that $d{\mathsf{P}}^{*}=(1/X^{*})d{\mathsf{Q}}$ yields an element of ${\mathcal{P}_{\textnormal{eff}}}$; and the first-order condition for $g({\mathsf{P}})$ states that $X^{*}=d{\mathsf{Q}}/d{\mathsf{P}}^{*}$ is an element of ${\mathcal{E}}$. Furthermore, the two objective functions are related by $f(X)\leq 0\leq g({\mathsf{P}})$, with equality precisely at the optimal solutions $X^{*}$ and ${\mathsf{P}}^{*}$. We refer to Ekeland and Témam (1999) for more information on convex optimization in infinite dimension; see in particular Chapter II therein.
@@
Following Remark 2.5, let us consider a benchmark $X_{\text{b}}$. In the presence of a benchmark, the strong duality relation (3.3) takes the form
${\mathbb{E}}_{\mathsf{Q}}\left[\log\frac{X^{*}}{X_{\text{b}}}\right]=\sup_{X\in{\mathcal{E}}}{\mathbb{E}}_{\mathsf{Q}}\left[\log\frac{X}{X_{\text{b}}}\right]=\inf_{{\mathsf{P}}\in{\mathcal{P}}_{\text{eff}}}{\mathbb{E}}_{\mathsf{Q}}\left[-\log\left(\frac{d{\mathsf{P}}^{a}}{d{\mathsf{Q}}}X_{\text{b}}\right)\right]={\mathbb{E}}_{\mathsf{Q}}\left[-\log\left(\frac{d{\mathsf{P}}^{*}}{d{\mathsf{Q}}}X_{\text{b}}\right)\right],$
where now all four quantities are finite. The optimizers $X^{*}$ and ${\mathsf{P}}^{*}$ do not depend on the choice of benchmark. This ‘benchmarked’ version of strong duality is proved by replacing $X$ by $X/X_{\text{b}}$ and $d{\mathsf{P}}^{a}/d{\mathsf{Q}}$ by $X_{\text{b}}(d{\mathsf{P}}^{a}/d{\mathsf{Q}})$ in (3.4) and the subsequent manipulations.
Remark 3.9.
In the spirit of Lardy et al. (2024), one can define the maximum description gain of switching from ${\mathsf{P}}\in{\mathcal{P}_{\textnormal{eff}}}$ to some other element of ${\mathcal{P}_{\textnormal{eff}}}$ by
$H({\mathsf{Q}}\mid{\mathsf{P}}\leadsto{\mathcal{P}_{\textnormal{eff}}})=\sup_{{\mathsf{P}}^{\prime}\in{\mathcal{P}_{\textnormal{eff}}}}{\mathbb{E}}_{\mathsf{Q}}\left[\log\frac{d{\mathsf{P}}^{\prime a}}{d{\mathsf{P}}}\right]$
if ${\mathsf{Q}}$ is absolutely continuous with respect to ${\mathsf{P}}$, and +$\infty$+ ∞ otherwise. Thanks to condition (iii) of Theorem 3.6, assuming that ${\mathsf{Q}}\ll{\mathcal{P}}$, the RIPr is characterized as the unique element of ${\mathcal{P}_{\textnormal{eff}}}$ equivalent to ${\mathsf{Q}}$ such that $H({\mathsf{Q}}\mid{\mathsf{P}}^{*}\leadsto{\mathcal{P}_{\textnormal{eff}}})=0$, and this is the smallest possible value that the maximum description gain can take.
Remark 3.10.
The strong duality in (3.3) implies that there exists an e-variable with positive expected logarithm under ${\mathsf{Q}}$ if and only if ${\mathsf{Q}}\notin{\mathcal{P}_{\textnormal{eff}}}$. Therefore, thanks to Corollary 4.6 below, Theorem 3.6 extends a result by Zhang et al. (2024) showing that if ${\mathcal{P}}$ is finite, then such an e-variable exists if and only if ${\mathsf{Q}}$ is not in the convex hull of ${\mathcal{P}}$.
Remark 3.11.
The RIPr ${\mathsf{P}}^{*}$ is not a probability measure in general. Equivalently, it may happen that ${\mathbb{E}}_{\mathsf{Q}}[1/X^{*}]<1$. We will see this phenomenon in Sections 5.1 and 5.4, and other examples are given by Lardy et al. (2024). It is also worth noting that there may exist elements $\widetilde{\mathsf{P}}\in{\mathcal{P}_{\textnormal{eff}}}$ that are different from ${\mathsf{P}}^{*}$, but whose absolutely continuous part $\widetilde{\mathsf{P}}^{a}$ is equal to ${\mathsf{P}}^{*}$. We will see this in Section 5.4 with $\widetilde{\mathsf{P}}$ being the symmetrization of ${\mathsf{Q}}$. Other examples are obtained from Example 3.12 below.
@@
Consider a simple null hypothesis ${\mathcal{P}}=\{{\mathsf{P}}_{0}\}$ with ${\mathsf{Q}}\ll{\mathsf{P}}_{0}$ so that the numeraire $X^{*}$ is finite. In this case ${\mathsf{P}}^{*}={\mathsf{P}}_{0}^{a}$ and $X^{*}=1/(d{\mathsf{P}}^{*}/d{\mathsf{Q}})=d{\mathsf{Q}}/d{\mathsf{P}}_{0}$ up to ${\mathsf{Q}}$-nullsets. Here one can easily check that ${\mathcal{P}_{\textnormal{eff}}}$ consists of all sub-probability measures ${\mathsf{P}}$ such that ${\mathsf{P}}(A)\leq{\mathsf{P}}_{0}(A)$ for all $A$. Indeed, any such ${\mathsf{P}}$ must belong to ${\mathcal{P}_{\textnormal{eff}}}$ since ${\mathbb{E}}_{\mathsf{P}}[X]\leq{\mathbb{E}}_{{\mathsf{P}}_{0}}[X]\leq 1$ for all e-variables $X$. Conversely, if ${\mathsf{P}}$ is not of this form, then ${\mathsf{P}}(A)>{\mathsf{P}}_{0}(A)$ for some $A$, and $X=\bm{1}_{A}/{\mathsf{P}}_{0}(A)$ is an e-variable such that ${\mathbb{E}}_{\mathsf{P}}[X]>1$, showing that ${\mathsf{P}}\notin{\mathcal{P}_{\textnormal{eff}}}$. If ${\mathsf{P}}_{0}(A)=0$ we set $X=\infty\bm{1}_{A}$ in the previous sentence.
We end this section with the general version of Theorem 3.6. The proof reduces the general case to the one treated in Theorem 3.6 by introducing the conditional probability measure ${\mathsf{Q}}^{*}(A)={\mathsf{Q}}(A\mid X^{*}<\infty)$ if the quantity $\lambda^{*}={\mathsf{Q}}(X^{*}<\infty)$ is positive, where $X^{*}$ is the numeraire. The fully degenerate case $\lambda^{*}=0$ is treated separately. It is evident from Definition 3.4 that the RIPr ${\mathsf{P}}^{*}$ is absolutely continuous with respect to ${\mathsf{Q}}$, and we know from Lemma 3.3 that it belongs to ${\mathcal{P}_{\textnormal{eff}}}$. As shown below, one actually has more precise information, namely that
@@
where we write $\lambda^{*}\cdot{\mathcal{P}_{\textnormal{eff}}}$ for the set $\{\lambda^{*}{\mathsf{P}}\colon{\mathsf{P}}\in{\mathcal{P}_{\textnormal{eff}}}\}\subset{\mathcal{P}_{\textnormal{eff}}}$. We continue to write ${\mathsf{P}}^{a}$ for the absolutely continuous part of ${\mathsf{P}}$ with respect to ${\mathsf{Q}}$. When applying this theorem it is useful to know that the event $\{X^{*}<\infty\}$ can be determined without first computing the numeraire; see Remark 3.14 below.
@@
    ${\mathbb{E}}_{\mathsf{Q}}\left[\frac{d{\mathsf{P}}^{a}}{d{\mathsf{P}}^{*}}\bm{1}_{\{X^{*}<\infty\}}\right]\leq 1\text{ for all }{\mathsf{P}}\in{\mathcal{P}_{\textnormal{eff}}}.$
@@
    ${\mathbb{E}}_{\mathsf{Q}}\left[\log\left(\lambda^{*}\frac{d{\mathsf{P}}^{a}}{d{\mathsf{P}}^{*}}\right)\bm{1}_{\{X^{*}<\infty\}}\right]\leq 0\text{ for all }{\mathsf{P}}\in{\mathcal{P}_{\textnormal{eff}}}.$
@@
${\mathbb{E}}_{{\mathsf{Q}}^{*}}[\log X^{*}]=\sup_{X\in{\mathcal{E}}}{\mathbb{E}}_{{\mathsf{Q}}^{*}}[\log X]=\inf_{{\mathsf{P}}\in{\mathcal{P}_{\textnormal{eff}}}}H({\mathsf{Q}}^{*}\mid{\mathsf{P}})=H({\mathsf{Q}}^{*}\mid{\mathsf{P}}^{**}).$ (3.7)
@@
We first let ${\mathsf{P}}^{*}$ be the RIPr and check (3.6). Equivalence with ${\mathsf{Q}}$ on $\{X^{*}<\infty\}$ is evident from Definition 3.4. To show that ${\mathsf{P}}^{*}$ belongs to $\lambda^{*}\cdot{\mathcal{P}_{\textnormal{eff}}}$ we refine the calculation in the proof of Lemma 3.3. Consider any $X\in{\mathcal{E}}$ and set $X^{\prime}=X\bm{1}_{\{X^{*}<\infty\}}+\infty\bm{1}_{\{X^{*}=\infty\}}$. This still belongs to ${\mathcal{E}}$ since it is ${\mathsf{P}}$-almost surely equal to $X$ for all ${\mathsf{P}}\in{\mathcal{P}}$. Thanks to the convention (1.1) we have $X^{\prime}/X^{*}=\bm{1}_{\{X^{*}<\infty\}}X^{\prime}/X^{*}+\bm{1}_{\{X^{*}=\infty\}}$ and hence
@@
using the numeraire property in the last step. It follows that ${\mathsf{P}}^{*}\in\lambda^{*}\cdot{\mathcal{P}_{\textnormal{eff}}}$.
@@
Since ${\mathsf{Q}}^{*}(X^{*}<\infty)=1$ by construction, we have from Theorem 2.6 that ${\mathsf{Q}}^{*}\ll{\mathcal{P}}$. We also define ${\mathsf{P}}^{**}={\mathsf{P}}^{*}/\lambda^{*}$ which, thanks to (3.6), is an element of ${\mathcal{P}_{\textnormal{eff}}}$ equivalent to ${\mathsf{Q}}^{*}$. Now, the statement (i) that ${\mathsf{P}}^{*}$ is the RIPr of ${\mathsf{Q}}$ is equivalent to ${\mathsf{P}}^{**}$ being the RIPr of ${\mathsf{Q}}^{*}$. Furthermore, (ii) and (iii) can be equivalently expressed in terms of ${\mathsf{Q}}^{*}$ and ${\mathsf{P}}^{**}$ as
${\mathbb{E}}_{{\mathsf{Q}}^{*}}\left[\frac{d{\mathsf{P}}^{a}}{d{\mathsf{P}}^{**}}\right]\leq 1\text{ for all }{\mathsf{P}}\in{\mathcal{P}_{\textnormal{eff}}}\quad\text{and}\quad{\mathbb{E}}_{{\mathsf{Q}}^{*}}\left[\log\frac{d{\mathsf{P}}^{a}}{d{\mathsf{P}}^{**}}\right]\leq 0\text{ for all }{\mathsf{P}}\in{\mathcal{P}_{\textnormal{eff}}},$
@@
Remark 3.14.
The event $\{X^{*}<\infty\}$ can be determined without first computing the numeraire. Let ${\mathsf{Q}}={\mathsf{Q}}^{r}+{\mathsf{Q}}^{s}$ be the generalized Lebesgue decomposition of ${\mathsf{Q}}$ with respect to ${\mathcal{P}}$ given by Lemma A.1, and let $N$ be the event given in the lemma. Since $N$ is ${\mathcal{P}}$-negligible, $X=\infty\bm{1}_{N}$ is an e-variable, so the numeraire must be infinite on $N$, ${\mathsf{Q}}$-almost surely. On the other hand, because the event $N^{\prime}=\{X^{*}=\infty\}$ is ${\mathcal{P}}$-negligible, the lemma also yields ${\mathsf{Q}}(N^{\prime}\setminus N)=0$. Thus $N$ and $\{X^{*}=\infty\}$ are both ${\mathcal{P}}$-negligible and coincide up to ${\mathsf{Q}}$-nullsets, so the former can be used in place of the latter in Theorem 3.13.
@@
Remark 3.14 implies that $N$ equals the event where the numeraire is infinite up to ${\mathsf{Q}}$-nullsets, so ${\mathsf{Q}}^{*}$ defined in the corollary coincides with ${\mathsf{Q}}^{*}$ defined in Theorem 3.13. Moreover, thanks to Example 3.12, the numeraire for ${\mathsf{Q}}^{*}$ is $d{\mathsf{Q}}^{*}/d{\mathsf{P}}_{0}$ up to ${\mathsf{Q}}^{*}$-nullsets. From these facts and Theorem 3.13, we conclude that the numeraire for ${\mathsf{Q}}$ is given by (3.8). Finally, since $1/X^{*}=d{\mathsf{P}}_{0}^{a}/d{\mathsf{Q}}$ up to ${\mathsf{Q}}$-nullsets, the RIPr is ${\mathsf{P}}^{*}={\mathsf{P}}_{0}^{a}$ as claimed. ∎
4 Finding the numeraire and RIPr
So far we have studied the abstract existence, uniqueness, and duality theory of the numeraire and RIPr, but we have not said much about how to actually compute them. In this section we develop tools for doing so, which we then apply to several examples in Section 5.
We focus on the statistically most relevant situation where ${\mathsf{Q}}\ll{\mathcal{P}}$. In this case the numeraire $X^{*}$, and indeed every e-variable, is ${\mathsf{Q}}$-almost surely finite and the RIPr ${\mathsf{P}}^{*}$ is equivalent to ${\mathsf{Q}}$. The general case can be reduced to this case by replacing ${\mathsf{Q}}$ with the conditional measure ${\mathsf{Q}}^{*}$ in Theorem 3.13 (except in the fully degenerate situation where the numeraire is infinity and the RIPr zero.)
Proposition 2.4 suggests that $X^{*}$ can be found by solving an optimization problem. Unfortunately this may not be straightforward because the optimization is over the set of all e-variables, which can be difficult to characterize. For the same reason, even if a candidate $X^{*}$ has been found it may not be clear how to check that it does, in fact, satisfy the numeraire property. Alternatively, the strong duality of Theorem 3.6 suggests minimizing relative entropy to find ${\mathsf{P}}^{*}$. However, the optimization is again over a set that can be difficult to describe explicitly, in this case the effective null ${\mathcal{P}_{\textnormal{eff}}}$. Moreover, checking that a candidate ${\mathsf{P}}^{*}$ is optimal involves comparing it to all elements of ${\mathcal{P}_{\textnormal{eff}}}$.
We first present a verification theorem which simplifies the task of checking that candidates $X^{*}$ and ${\mathsf{P}}^{*}$ are in fact optimal. This result strengthens Corollary 2 of Grünwald et al. (2024). Despite the simple proof, the result turns out to be remarkably useful.